2. Find : (i) 93/2 (ii) 322/5 (iii) 163/4 (iv) 125-1/3

2. Find :
(i) 93/2 (ii) 322/5
(iii) 163/4 (iv) 125-1/3

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 93/2 = (32)3/2 = (3)2#3/2 = (3)3 = 27 (ii) 322/5 = (25)2/5 = 25#2/5 = (2)2 = 4 (iii) 163/4 = (24)3/4 = (2)4#3/4 = (2)3 = 8 (iv) 125-1/3 = (53)−1/3 = (5)3 #(−1/3) = (5)−1 = 15